Lisa can shovel snow from her driveway in 65 minutes. Jim can do the same job in 55 minutes. How long would it take Lisa and Jim to shovel the driveway together?
Answer by checkley77(12844)

1/65+1/55=1/X
(55+65)/55*65=1/X
120/3,575=1/X
120X=3,575
X=3,575/120
X=29.79 MINUTES.